OpenStudy (animalover02):

I WILL MEDAL & FAN!!!!!!!!! What is the solution to the system of equations? 6x-2y=2 -3x+4y=5 A. (1 , -2) B. (2 , 3) C. (1 , 3) D. ( 1 , 2)

OpenStudy (animalover02):

@ilianapaulk

OpenStudy (dopeboydee):

It is D.

OpenStudy (animalover02):

@ilianapaulk do you agree with @DopeBoyDee

OpenStudy (animalover02):

i thought D but wasnt sure

OpenStudy (mathmale):

Kindly do not guess. Please show the work you did to justify your choice of answer.

OpenStudy (mathmale):

"thinking it's D" doesn't make it so.

OpenStudy (animalover02):

ik that..id like help not doing it by myself. these are hard to thats why i need help

OpenStudy (ilianapaulk):

Solve equation [2] for the variable x Plug this in for variable x in equation [1] Solve equation [1] for the variable y

OpenStudy (dopeboydee):

yea ^

OpenStudy (mathmale):

Do what you can and share that. Then others can see where you need help and jump in with appropriate comments. Look at ilianapaulk"s suggestion, a very good one. You are responsible for learning how to do that.

OpenStudy (animalover02):

i solved it out on paper and it came out D. i just want to know if im correct. i dont have much time left:)

OpenStudy (mathmale):

You want to eliminate one variable (x or y) to find the value of the other variable. You are given 6x-2y=2 -3x+4y=5 I'd suggest you multiply the 2nd equation by 2, all the way through. Write your result here:

OpenStudy (animalover02):

(1 , 2)

OpenStudy (dopeboydee):

its not C

OpenStudy (animalover02):

im going with D any of youll against the answer or is it corrct like i said???

OpenStudy (dopeboydee):

its correct i did the problem to

OpenStudy (mathmale):

6x-2y=2 -3x+4y=5 I asked you to multiply the 2nd equation by 2. Result: -6x+8y=10. Now add the first and 3rd equations together. Show your work.

OpenStudy (mathmale):

If you want to go with D, then substitute your x and y values into both of the given equations. Are both equations then true? If so, D is your solution. You can check your work yourself by doing this.

OpenStudy (animalover02):

thanks everyone!

OpenStudy (dopeboydee):

welcome :)

OpenStudy (ilianapaulk):

i agree had to work out the problem myself ... just take your time on questions like those cause they can be tricky and confusing if you miss one step D is right